# Perfect Matching in Product Graphs and in their Random Subgraphs

### v2

@@ 3.3 Typically no obstructions
We follow a similar argument to that in Lemma 3.12. Fix $u,w,s$, and $b$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $\frac{n}{2}\neq b=|B|$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
By 1.2 $e(B,B^{C})\geq f(b)=f(n-b)$ and we obtain
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+f(n-b)+wd-w.$
By the proof of Lemma 3.3, there are at most $n^{\ell_{2}+w/2}(ed)^{2s+w/2}$ ways to choose $W\cup S$. Furthermore, since we assume $e(B,B^{C})\leq|B^{C}|\ln^{2}d$ and $|B^{C}|\geq u\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2|B^{C}|}{\ln d}\right\}=\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ choices for $B^{C}$, and hence for $B$. Therefore, by Lemma 3.2, there are at most $2n^{\ell_{2}+w/2}(ed)^{2s+w/2}\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ such minimal obstructions. We thus obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(n-b)}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(n-b)}$
$\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell_{1}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and by Claim 3.6 $\ell_{3}\leq\frac{6u}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell-\ell_{2}-\ell_{3}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}$
$\leq 2\exp\left\{\left(\frac{6u}{d^{2}}+\epsilon\frac{s}{3}-s(1-2\epsilon)-(1-\epsilon)(\ell-u)\right)\ln n+\frac{2(n-b)}{\ln d}-pf(n-b)\right\}$
$\leq 2\exp\left\{\frac{3(n-b)}{\ln d}-pf(n-b)\right\},$
where in the last step we used that $u<\ell$. Since $b\geq\frac{n}{2}$, we have that $\log_{C}(n-b)\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f(n-b)\geq\frac{n-b}{C^{2}}$. Moreover, $n-b\geq u$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{u\left(\frac{3}{\ln d}-\frac{1}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{u}{2C^{2}}\right\}.$
A union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
@@ 4 Discussion
Let $G$ be a $d$-regular graph on $n$ vertices, with $d=\omega(1)$ and $n$ divisible by two. What minimal requirements are needed on $G$, such that in the graph process on $G$, the hitting times for minimum degree one, connectivity, and the existence of a perfect matching are the same?
@@ Acknowledgements
The authors would like to thank Joshua Erde, Mihyun Kang and Michael Krivelevich for their guidance, advice, and fruitful discussions. The second author was supported in part by the Austrian Science Fund (FWF) [10.55776/W1230].

### v3

@@ 3.3 Proof of Theorem 3b
Similarly to before, we have that
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+f^{*}(n-b).$
Since we assume $e(B,B^{C})\leq|B^{C}|\ln^{2}d$ and $|B^{C}|\geq u\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2|B^{C}|}{\ln d}\right\}=\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ choices for $B^{C}$, and hence for $B$. Therefore, by Lemma 3.3, there are at most $2\exp\left\{\frac{2(n-b)}{\ln d}\right\}$ such minimal obstructions. We thus obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{(\hat{e}-du)/2}$
$\leq 2\exp\left\{\frac{2(n-b)}{\ln d}\right\}(1-p)^{\frac{1}{2}d(\ell_{1}-u)+\frac{f^{*}(n-b)}{2}}$
$\leq 2\exp\left\{\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}\right\}\left(\frac{\omega}{n}\right)^{\frac{1}{2}(\ell_{1}-u)}$
$\leq 2\exp\left\{-\frac{1}{2}(\ell_{1}-u)\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}.$
Recall that $\ell=\ell_{1}+\ell_{3}$ and by Claim 3.18 $\ell_{3}\leq\frac{9u}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{-\frac{1}{2}(\ell-u)\ln n+\frac{1}{2}\ell_{3}\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{9u}{2d^{2}}\ln n+\frac{2(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{3(n-b)}{\ln d}-\frac{pf^{*}(n-b)}{2}\right\},$
where in the penultimate step we used that $u<\ell$ and in the last step we used that $\frac{u}{d^{2}}\ln n\ll\frac{n-b}{\ln d}$ and that $\omega$ tends to infinity arbitrarily slowly. Since $b\geq\frac{n}{2}$, we have that $\log_{C}(n-b)\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f^{*}(n-b)\geq\frac{n-b}{C^{2}}$. Moreover, $n-b\geq u$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{u\left(\frac{3}{\ln d}-\frac{1}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{u}{2C^{2}}\right\}.$
Recalling that $u\geq\frac{n}{d^{C^{5}/p}}$ a union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
We can now conclude our proof.
Proof of Theorem 3b.
By Lemmas 3.10 and 3.15, we have that whp there are no minimal non-trivial obstructions with $u\leq\frac{n}{d^{C^{5}/p}}$. By Lemmas 3.16 and 3.17, we have shown that whp there are no minimal non-trivial obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$ and $b\leq n/2$. By Lemmas 3.19 and 3.20, we have shown that whp there are no minimal non-trivial obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$ and $b\geq n/2$. Thus, we have shown that whp there are no minimal non-trivial obstructions. Thus, by the Tutte-Berge formula, it follows that the giant component in $G_{p}$ has a (nearly-)perfect matching. ∎
@@ 4 Discussion
Let $G$ be a $d$-regular graph on $n$ vertices, with $d=\omega(1)$ and $n$ divisible by two. What minimal requirements are needed on $G$, such that in the random graph process on $G$, the hitting times for minimum degree one, connectivity, and the existence of a perfect matching are the same?
@@ Acknowledgements
The authors would like to thank Joshua Erde, Mihyun Kang and Michael Krivelevich for their guidance, advice, and fruitful discussions. Special thanks to an anonymous referee for spotting an error in a previous version. The second author was supported in part by the Austrian Science Fund (FWF) [10.55776/W1230].